OpenStudy (anonymous):

PLEASE HELP! write a quadratic equation that has the given roots 6,-9

OpenStudy (lucaz):

the roots are the values that make the equation equal to zero

OpenStudy (lucaz):

you're given 6 and -9, so f(6) and f(-9) = 0

OpenStudy (lucaz):

you can write f(x)=(x-6)(x+9)

OpenStudy (ybarrap):

Roots mean $$ f(x)=(x-x_1)(x-x_2)=0\\ $$ where \(x_1\) and \(x_2\) are the roots.

OpenStudy (anonymous):

then factor?

OpenStudy (anonymous):

foil i mean

OpenStudy (anonymous):

not factor lol

OpenStudy (lucaz):

this form is already factored

OpenStudy (anonymous):

i mean foil to get full quadratic equation

OpenStudy (lucaz):

foil gives you the standard form \[f(x)=ax^2+bx+c\]

OpenStudy (anonymous):

yes thats what i need thank you

OpenStudy (lucaz):

no problem